The area of a rectangle is 360 inches^2. The ratio of the length to the width is 5:2. Find the length and the width
Answer by lwsshak3(11628)

The area of a rectangle is 360 inches^2. The ratio of the length to the width is 5:2. Find the length and the width
***
let x=length
let y=width
x/y=5/2
y=2x/5
area=length*width
area=x*y=x(2x/5)=2x^2/5
2x^2/5=360
2x^2=1800
x^2=900
x=30
y=60/5=12
length=30 inches
width=12 inches